OpenStudy (anonymous):

The length of a diagnol square is 7.5 mm. FInd the area of the square

OpenStudy (anonymous):

Use pythagorean theorem. If it is a square, then all sides must be equal in length. Call this length \[x\]

OpenStudy (anonymous):

The diagonal is then \[\sqrt{x^2+x^2}=\sqrt{2x^2}\]

OpenStudy (anonymous):

Since this is equal to 7.5mm, \[\sqrt{2x^2}=7.5\]

OpenStudy (anonymous):

Solve for x, which represents the length of a side of the square.

OpenStudy (anonymous):

Since area is length times width, the area is just \[x(x)\]

OpenStudy (anonymous):

=x^2